# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v2

@@ D.3.2 Stable invariants
where we take $\rho,\sigma\in\mathcal{R}$, $N=\mathrm{dim}(\rho)$, and $M=\mathrm{dim}(\sigma)$. The stabilized space $\mathrm{QCA}^{\mathcal{R}}_{0}$ is the direct limit
@@
$\tilde{W}^{\rho}_{n}[V]=\frac{W^{\rho}_{n}[V]}{\rho}\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))],$ (140)
which then obeys $\tilde{W}^{\rho\sigma}_{n}[V\otimes\mathbbm{1}]=\tilde{W}^{\rho}_{n}[V]$. Here, $\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))]$ is the algebraic localization of the module $\mathrm{Rep}(G)[\pi_{n}(U(N))]$, defined in the next section.
@@
$\tilde{w}^{\rho}_{1}[V]=\tilde{W}^{\rho}_{1}[V]+1\cdot{\mathbb{Z}}\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{1}(U(N))]/(1\cdot{\mathbb{Z}}).$ (141)
D.3.3 Classification
@@
The formal ratios $\tilde{W}^{\rho}_{n}[V]$ are properly defined as elements of the localization $\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))]$, which from Eqs. (138143) gives
$\tilde{W}^{\rho\sigma}_{n}[V\otimes Q]=\tilde{W}^{\rho}_{n}[V]+\tilde{W}^{\sigma}_{n}[Q].$ (145)
@@
$\tilde{w}^{\rho}_{1}[V]\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{1}(U(N))]/(1\cdot\pi_{1}(U(N))).$ (146)
@@
Note that Eq. (142) implies that $\tilde{W}^{\rho_{1}}_{n}[V]=\tilde{W}^{\rho_{2}}_{n}[Q]$ if and only if there are representations $\sigma\otimes\rho_{1}$ and $\sigma\otimes\rho_{2}$ in $\mathcal{R}$ such that the unstable winding numbers
@@
D.4 One dimension with symmetry
The study of the connected components $\pi_{0}(\mathrm{QCA}_{1}^{\mathcal{R}})$ has been considered in Refs. [93, 26, 114], where complete classifications were found for $U(1)$ and finite cyclic groups of prime order. In this appendix, we use the representation theory technology of the previous section to generalize these classifications to arbitrary topological groups $G$ and countable collections of ancillae representations $\mathcal{R}$.
D.4.1 Unstable invariants
where
$u_{x,x+1}\in U(N_{x}N_{x+1}\to N^{\prime}_{x}N^{\prime}_{x+1}).$ (154)
Then $V$ commutes with the on-site symmetry $\prod_{x}\rho_{x}$ if and only if
$u_{x,x+1}\rho_{x}(g)\rho_{x+1}(g)=v_{x}(g)v_{x+1}(g)u_{x,x+1}$ (155)
The product of two consecutive $v_{x}v_{x+1}$ is, by Eq. (155), unitarily equivalent to the representation $\rho_{x}\rho_{x+1}$. However, this only implies that $v_{x}$ and $v_{x+1}$ are projective representations of $G$. The classifications of Ref. [93, 26, 114] rely on the following observation: $V$ is equal to a finite depth circuit with symmetric unitary gates (which implies they are connected by a symmetric path) if and only if there are functions $\phi_{x}:G\to U(1)$ such that $e^{i\phi_{x}}v_{x}$ is unitarily equivalent to $\rho_{x}$ for all $x$. That is, the projective representation $v_{x}$ can be lifted to an ordinary unitary representation, and that representation is precisely $\rho_{x}$. A more concise statement of the condition is that $v_{x}$ and $\rho_{x}$ are isomorphic when both regarded as projective representations.
Indeed, suppose that
with
such that
$u^{\prime}_{x,x+1}\rho_{x}\rho_{x+1}$ $=w_{x}w_{x+1}v_{x}v_{x+1}u_{x,x+1}$ (159)
×$\qquad\qquad\qquad\times w_{x}w_{x+1}u_{x,x+1}$ (160)
$=\rho_{x}\rho_{x+1}u^{\prime}_{x,x+1}$ (161)
is symmetric.
Conversely, if $V$ can be written as a circuit with symmetric unitary gates, as in Eq. (157), then we must have that
$\left(\prod_{x\in 2{\mathbb{Z}}+1}u_{x,x+1}u^{\prime\dagger}_{x,x+1}\right)\left(\prod_{x\in 2{\mathbb{Z}}}u^{\prime}_{x,x+1}u^{\dagger}_{x,x+1}\right)=\mathbbm{1}.$ (162)
$u^{\prime}_{x,x+1}u^{\dagger}_{x,x+1}=w_{x}w_{x+1}$ (163)
must be a tensor product of isometries $w_{x}\in U(N_{x}\to N_{x}^{\prime})$. Further, as the whole circuit is the identity, we have $w_{x}^{\dagger}w_{x}=\mathbbm{1}$ and thus $N_{x}^{\prime}\geq N_{x}$ for all $x$. However, $N_{x}N_{x+1}=N^{\prime}_{x}N^{\prime}_{x+1}$, and these conditions can only be simultaneously satisfied if $N_{x}=N^{\prime}_{x}$ for all $x$, and thus $w_{x}$ is a unitary gate.
Then we have
$v_{x}v_{x+1}u_{x,x+1}$ $=u_{x,x+1}\rho_{x}\rho_{x+1}$ (164)
$=w_{x}w_{x+1}u^{\prime}_{x,x+1}\rho_{x}\rho_{x+1}$ (165)
$=w_{x}w_{x+1}\rho_{x}\rho_{x+1}u^{\prime}_{x,x+1}$ (166)
so that $v_{x}=e^{-i\phi_{x}}w_{x}\rho_{x}w^{\dagger}_{x}$, as we needed.
Thus, an invariant for symmetric QCA in one dimension must somehow measure the difference between $v_{x}$ and $\rho_{x}$. We split this into two parts. First, we need $v_{x}$ to lift to a usual unitary representation. It is well known that the obstruction to lifting a projective unitary representation to a unitary representation is captured by group cohomology with coefficients in $U(1)$, so there is an
$\omega_{x}\in H^{2}(G;U(1))$ (168)
such that $v_{x}$ lifts to a unitary representation if and only if $\omega_{x}=0$. (We use an additive notation for $\omega_{x}$.) Further, $e^{-i\omega_{x}}v_{x}$ always lifts to a unitary representation, say $\tilde{v}_{x}$. The unitary representation $\tilde{v}_{x}$ is unique up to multiplication by a one dimensional representation $\chi_{x}:G\to U(1)$. Thus, we can identify the isomorphism class of $\tilde{v}_{x}$, with a coset in the representation rig,
$\tilde{v}_{x}X_{G}\in\mathrm{Rep}(G)/X_{G},$ (169)
where $X_{G}$ is the multiplicative subgroup of units (one dimensional representations) in $\mathrm{Rep}(G)$. Then $V$ is a symmetric circuit if and only if
$(\omega_{x},\tilde{v}_{x}X_{G})=(0,\rho_{x}X_{G})$ (170)
for all $x$.
D.4.2 Stable invariants
While the $\omega_{x}$ invariant is stable to stacking ancillae, the $\tilde{v}_{x}X_{G}$ invariant is not, and both apparently depend on position $x$. Stabilizing the $\tilde{v}_{x}X_{G}$ invariant through the algebraic localization construction fixes both of these issues.
We modify the $\tilde{v}_{x}X_{G}$ invariant to
$\frac{v_{x}}{\rho_{x}}X_{G}\in\mathcal{R}^{-1}\mathrm{Rep}(G)/X_{G},$ (171)
which we now regard as a multiplicative group via the multiplication in $\mathrm{Rep}(G)$.
The quotient by $X_{G}$ may refer to either the quotient by $X_{G}/1\subseteq\mathcal{R}^{-1}\mathrm{Rep}(G)$ or performing the $X_{G}$ quotient first, then the localization. The results are isomorphic.
When appending additional ancillae to site $x$ with representation $\sigma_{x}$, $V\otimes\mathbbm{1}$ acts trivially on the ancilla, so $v_{x}$ just becomes $v_{x}\otimes\sigma_{x}$. Then $v_{x}\sigma_{x}/\rho_{x}\sigma_{x}=v_{x}/\rho_{x}$ is unchanged.
Further, because $v_{x}v_{x+1}=\rho_{x}\rho_{x+1}\in\mathcal{R}$, we have
$\omega_{x}+\omega_{x+1}=0\quad\text{and}\quad\frac{\tilde{v}_{x}}{\rho_{x}}\frac{\tilde{v}_{x+1}}{\rho_{x+1}}=1.$ (172)
So the invariant for $x+1$ is just the inverse of that for $x$, and has no new information. Thus, we can just focus on $x=0$, and henceforth drop the $x$ index.
D.4.3 Classification
We have a complete set of invariants in $H^{2}(G;U(1))\times\mathcal{R}^{-1}\mathrm{Rep}(G)/X_{G}$. Again, not all $(\omega,\tilde{v}X_{G}/\rho)$ can actually be achieved with a fixed $\mathcal{R}$. We focus primarily on $\tilde{v}X_{G}/\rho$.
The only numerators $\tilde{v}$ which may occur in the invariant $\tilde{v}X_{G}/\rho$ are those which occur as tensor factors in $\rho$ for some $\rho\in\mathcal{R}$. That is, the numerators belong to the saturation
^=$\hat{\mathcal{R}}=\{v\in\mathrm{Rep}(G)\,:\,\exists\rho\in\mathrm{Rep}(G)\\ \text{such that }v\otimes\rho\in\mathcal{R}\}.$ (173)
Note that $\hat{\mathcal{R}}$ is closed under multiplication by $\mathcal{R}$, so the localization $\mathcal{R}^{-1}\hat{\mathcal{R}}$ is well defined and a multiplicative group under the natural multiplication operation of fractions.
All the elements of $\hat{\mathcal{R}}$ may occur as invariants. Suppose $v_{0}/\rho_{0}\in\mathcal{R}^{-1}\hat{\mathcal{R}}$. Then, as $v_{0}$ belongs to $\hat{\mathcal{R}}$, there are representations $v_{1}\in\mathrm{Rep}(G)$ and $\rho_{1}\in\mathcal{R}$ and a unitary $u_{0,1}^{\prime}$ such that
$v_{0}\otimes v_{1}u_{0,1}^{\prime}=u_{0,1}^{\prime}\rho_{1}.$ (174)
We tensor in $\rho_{0}$ to get
where $u_{x,x+1}$ is a translation of $u_{0,1}$ from Eq. (175). The invariant for the QCA $V$ is, by construction, $(v_{0}X_{G})/\rho_{0}$, or its inverse $(v_{1}\rho_{0}X_{G})/\rho_{1}$.
With regard to the invariant described by Eq. (168), let us briefly address whether all elements of $H^{2}(G,U(1))$ can actually be achieved by a QCA [recall that this was not necessarily the case for the $H^{1}(G,U(1))$ invariant discussed in $d=0$, for infinite groups $G$]. This invariant characterizes the obstruction in lifting the projective representation $v_{x}$ to a unitary representation. For finite groups, all such invariants can be achieved. Taking the local Hilbert space to be the span of $|g\rangle$ with $g\in G$, and given any 2-coycle $\alpha:G\times G\to U(1)$, we define
$v_{0}(g)=\rho_{\mathrm{reg}}(g)D(g)$ (177)
where $\rho_{\mathrm{reg}}$ is the regular representation and $D(g)$ is the diagonal matrix with elements $\langle h|D(g)|h\rangle=\alpha(g,h)$. Then $v_{0}$ is known to be a projective representation with obstruction $[\alpha]\in H^{2}(G;U(1))$. By a similar construction to the previous paragraph, we can construct a QCA (indeed, a finite depth circuit) from the unitary gate $u_{0,1}$ which conjugate $\rho_{\mathrm{reg}}\otimes\rho_{\mathrm{reg}}$ to $v_{0}\otimes v_{1}$, where $v_{1}$ has obstruction $-[\alpha]$. This gate exists by the vanishing of the obstruction for $v_{0}\otimes v_{1}$. We leave open the precise restrictions on $\omega\in H^{2}(G;U(1))$ imposed by limited ancillae representations or by $G$ being infinite.
The full classifying group is then a subgroup
$\pi_{0}(\mathrm{QCA}_{1}^{\mathcal{R}})\leq H^{2}(G;U(1))\oplus\log\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{G},$ (178)
with only the first summand possibly allowing for this to be a proper subgroup. That is, the projection to the second summand is surjective. We express the second summand with a logarithm just to make the group operation addition. The induced homomorphism from evaluation on a product state, $e_{0}^{*}$, just corresponds to projecting to the first summand, and thus reproduces the group cohomology classification of invertible states in one dimension [66, 93].
The case of $G=U(1)$ again serves as a simple example [26, 27]. We take $\mathcal{R}\cong{\mathbb{N}}[x,x^{-1}]$ to consist of all representations of $U(1)$. Then ^$\hat{\mathcal{R}}\cong{\mathbb{N}}[x,x^{-1}]$ as well, and the algebraic localization is isomorphic to the multiplicative group of rational polynomials with positive coefficients,
$\mathcal{R}^{-1}\hat{\mathcal{R}}\cong\mathrm{Frac}\,{\mathbb{N}}[x]^{\times}.$ (179)
The group of units $X_{U(1)}$ consists of monomials, $x^{m}$. Expressing the quotient in the additive notation, this is
$\log\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{U(1)}\cong\log\mathrm{Frac}\,{\mathbb{N}}[x]^{\times}/{\mathbb{Z}}\log x.$ (180)
As $H^{2}(U(1);U(1))\cong 0$ is trivial, this is the complete classification of $U(1)$-symmetric QCA in one dimension.
It has also been observed that including the regular representation for a finite group $G$ in $\mathcal{R}$, $\rho_{\mathrm{reg}}\in\mathcal{R}$, reduces the classification to just $H^{2}(G;U(1))\oplus\log{\mathbb{Q}}^{\times}$ [93]. Indeed, in the current presentation it is clear that this is because $R_{G}$ has zero divisors when $G$ is a finite group. If $\rho_{\mathrm{reg}}$ is the regular representation, then in the representation rig we have $\rho_{\mathrm{reg}}\sigma=\mathrm{dim}(\sigma)\rho_{\mathrm{reg}}$, and so
$\rho_{\mathrm{reg}}[\sigma\mathrm{dim}(\sigma^{\prime})-\sigma^{\prime}\mathrm{dim}(\sigma)]=0,$ (181)
for any $\sigma,\sigma^{\prime}\in\mathrm{Rep}(G)$. Thus, $\sigma/\sigma^{\prime}=\mathrm{dim}(\sigma)/\mathrm{dim}(\sigma^{\prime})\in\mathcal{R}^{-1}\hat{\mathcal{R}}$, and any two representations of the same dimension define the same element of the algebraic localization, which is then isomorphic to ratios of positive integers, $\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{G}\cong{\mathbb{Q}}^{\times}$. (More correctly, this should itself be a localization of the natural numbers, $\mathrm{dim}(\mathcal{R})^{-1}\mathrm{dim}(\hat{\mathcal{R}})$.)
D.5 Functorality
@@
[…]
Further, we have that $(\rho_{1}\oplus\rho_{2})\phi\simeq\rho_{1}\phi\oplus\rho_{2}\phi$ and $(\rho_{1}\otimes\rho_{2})\phi\simeq\rho_{1}\phi\otimes\rho_{2}\phi$, so that $\rho\mapsto\rho\phi$ defines an induced rig homomorphism
@@
What’s more, any symmetric $V\in\mathrm{QCA}_{d}^{\mathcal{R}}$ (or more generally a QCA subject to some $G$ action, not necessarily symmetric) can also be regarded as being symmetric with respect to the $H$ action $\rho\phi$. Thus, we have a continuous inclusion map
@@
where ϕRep()subscriptsuperscriptitalic-$\phi^{*}_{\mathrm{Rep}}(\mathcal{R})\subseteq\mathcal{R}^{\prime}$. In turn, this continuous map induces homomorphisms
@@
In Appendices D.3 and D.4, we classified homotopy groups of symmetric QCA with $d-n\leq 1$. More concretely, we constructed isomorphisms between $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ and certain groups defined in terms of $\mathrm{Rep}(G)$. One should want these isomorphisms to be functorial—the induced map $\phi^{*}_{n}$ should be reproduced by applying $\phi^{*}_{\mathrm{Rep}}$ (or more correctly the induced map on the localization or quotient) to the invariant in $\mathcal{R}^{-1}\mathrm{Rep}(G)$.
@@
D.6 Lattice symmetries
It is believed that the generalized cohomology classification of invertible states also captures lattice symmetries and antiunitary symmetries [103, 104]. We also believe this to be true for QCA, and in this section we outline the structure that this should take. We leave detailed explorations to Ref. [107] and future work.
In the picture of Appendix C.3, a QCA on some $d$-dimensional orientable manifold $M$ can be identified by a map from $M$ to $\mathrm{QCA}_{d}$
$V\in\mathrm{QCA}_{M}\longleftrightarrow v:M\to\mathrm{QCA}_{d}.$ (189)
More precisely, we suspect $\mathrm{QCA}_{M}$ and $\left[M,\mathrm{QCA}_{d}\right]$ to be homotopy equivalent. Lattice symmetries can be treated in this picture as actions of the symmetry group $G$ on the base manifold $M$ [116]. Then, identifying the QCA $V$ with the map $v$, the QCA is symmetric if

### v3

@@ C.1.2 Stable invariants
where we take $\rho,\sigma\in\mathcal{R}$ (the monoid of ancillae), $N=\mathrm{dim}(\rho)$, and $M=\mathrm{dim}(\sigma)$. The stabilized space $\mathrm{QCA}^{\mathcal{R}}_{0}$ is the direct limit
@@
$\widetilde{W}^{\rho}_{n}[V]=\frac{W^{\rho}_{n}[V]}{\rho}\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))],$ (197)
which then obeys $\widetilde{W}^{\rho\sigma}_{n}[V\otimes\mathbbm{1}]=\widetilde{W}^{\rho}_{n}[V]$. Here, $\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))]$ is the algebraic localization of the module $\mathrm{Rep}(G)[\pi_{n}(U(N))]$, defined in the next section.
@@
$\widetilde{W}^{\rho}_{1}[V]=\widetilde{W}^{\rho}_{1}[V]+1\cdot{\mathbb{Z}}\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{1}(U(N))]/(1\cdot{\mathbb{Z}}).$ (198)
C.1.3 Classification
@@
The formal ratios $\widetilde{W}^{\rho}_{n}[V]$ are properly defined as elements of the localization $\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))]$, which from Eqs. (195200) gives
$\widetilde{W}^{\rho\sigma}_{n}[V\otimes Q]=\widetilde{W}^{\rho}_{n}[V]+\widetilde{W}^{\sigma}_{n}[Q].$ (202)
@@
$\widetilde{W}^{\rho}_{1}[V]\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{1}(U(N))]/(1\cdot\pi_{1}(U(N))).$ (203)
@@
Note that Eq. (199) implies that $\widetilde{W}^{\rho_{1}}_{n}[V]=\widetilde{W}^{\rho_{2}}_{n}[Q]$ if and only if there are representations $\sigma\otimes\rho_{1}$ and $\sigma\otimes\rho_{2}$ in $\mathcal{R}$ such that the unstable winding numbers
@@
C.2 Functorality
@@
$\rho\phi:H\xrightarrow{\phi}G\xrightarrow{\rho}U(N).$ (211)
Further, we have that $(\rho_{1}\oplus\rho_{2})\phi=\rho_{1}\phi\oplus\rho_{2}\phi$ and $(\rho_{1}\otimes\rho_{2})\phi=\rho_{1}\phi\otimes\rho_{2}\phi$, so that $\rho\mapsto\rho\phi$ defines an induced rig homomorphism
@@
What’s more, any symmetric $V\in\mathrm{QCA}_{d}^{\mathcal{R}}$ can also be regarded as being symmetric with respect to the $H$ action $\rho\phi$. Thus, we have a continuous inclusion map
@@
where $\phi^{*}_{\mathrm{Rep}}(\mathcal{R})\subseteq\mathcal{R}^{\prime}$. In turn, this continuous map induces homomorphisms
@@
In Appendix C.1 and Sec. IV.2, we classified homotopy groups of symmetric QCA with $d-n\leq 1$. More concretely, we constructed isomorphisms between $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ and certain groups defined in terms of $\mathrm{Rep}(G)$. One should want these isomorphisms to be functorial—the induced map $\phi^{*}_{n}$ should be reproduced by applying $\phi^{*}_{\mathrm{Rep}}$ (or more correctly the induced map on the localization or quotient) to the invariant in $\mathcal{R}^{-1}\mathrm{Rep}(G)$.
@@
C.3 Application to the QP pump
Finally, we briefly explain where the instance of the QP pump of Sec. V.3 sits in the classification of Appendix C.1. Specifically, we find its associated invariant.
The QP pump is $U(1)$-symmetric and has $d=1$, $n=2$. It thus has the same classifying group as loops of unitaries in $d=0$ with $U(1)$ symmetry. Indeed, this is the object from which our model of the QP pump was constructed. From Appendix C.1, to get the classification of $v(\theta)$ we must find its one-dimensional winding number in each symmetry sector. We have
Thus, up to a global phase, $v(\theta)$ has winding number $W$ in the charge 0 sector, and winding number 0 in the charge 1 sector.
$\widetilde{W}_{1}^{Q}[v]=\frac{1\cdot W+x\cdot 0}{1+x}+{\mathbb{Z}}=\frac{W}{1+x}+{\mathbb{Z}}.$ (218)
We can also see the functorality of Appendix C.2 in action. Suppose that we explicitly break the symmetry. Then the charge sector $1\leftrightarrow x$ should be mapped to having charge 0. The new invariant can be found by evaluating the rational function $\widetilde{W}_{1}^{Q}[v]$ at $x=1$. We find
$\widetilde{W}_{1}^{Q}[v]\mapsto\frac{W}{2}+{\mathbb{Z}}\in{\mathbb{Q}}/{\mathbb{Z}}.$ (219)
Now, all even $W$ are equivalent to the trivial phase. Interestingly, the odd $W$ phases remain non-trivial. They no longer support a quantized edge current, as we can deform the model with $W=+1$ to W=1W=-1, but there is still some nontrivial dynamics at the edge.
Indeed, we believe the odd $W$ phase without symmetry to be the two-tone-driven phase of Refs. [29, 30]. While the solvable limit of the model in Ref. [29] is highly singular, which makes computing invariants from integral formulae inconvenient, the edge phenomenology described in Ref. [29] is consistent with what we expect from having $\widetilde{W}_{1}\neq 0\bmod 1$. Namely, the edge dynamics was argued to be anomalous in that it could not be achieved by a zero-dimensional localized [in the sense that a continuous micromotion operator exists as in Eq. (133)] system which did not occur at the edge of an extended one-dimensional system. As the model in Ref. [29] is based around a version of the AKLT model, which can be trivialized when stacked with itself, it is thus reasonable to suppose that $\widetilde{W}_{1}=1/2\bmod 1$ for that model, as this is the only element of $\mathbb{Q}$$\mathbb{Z}$} of order 2. Thus, the abstract classification reveals a relationship between the interacting QP pump and Refs. [29, 30] which was not clear previously. It also extends the arguments for stability in Ref. [29] to the nonperturbative regime, where a picture of emergent symmetry (as employed in that reference) becomes inapplicable.
Appendix D QCA on manifolds or with spatial symmetries
In this appendix, we present conjectures regarding the classification of QCA on manifolds other than ${\mathbb{R}}^{d}$ (as considered in the main text), or with lattice symmetries.
The $\Omega$-spectrum property discussed in Sec. III.1 assumes that QCA are defined on the integer lattice ${\mathbb{Z}}^{d}$. By coarse graining, any locally finite set of points in ${\mathbb{R}}^{d}$ can be deformed to ${\mathbb{Z}}^{d}$, so the QCA is more properly based on the manifold ${\mathbb{R}}^{d}$. This distinction is important, as the classification of QCA based on the $d$-dimensional manifold $M$, $\mathrm{QCA}_{M}$, depends on the homology of $M$ [61] and not just its dimension.
$\mathrm{QCA}_{M}\to\left[M,\mathrm{QCA}_{d}\right]$ (220)
This leads to the hypothesis that, for orientable $M$ [120],
While we leave the general verification of this conjecture to future work, we remark here that Eq. (221) reproduces the classification of QCA on two-dimensional orientable manifolds in Ref. [61]. For non-orientable manifolds, one must consider twisted generalized cohomology—that is, $\mathrm{QCA}_{M}$ would be homotopy equivalent to spaces of sections in a $\mathrm{QCA}_{d}$-bundle over $M$ [64, 65].
The conjecture of Eq. (221) can be extended to include systems with lattice symmetries. Indeed, the generalized cohomology classification of invertible states is also believed to capture lattice symmetries and antiunitary symmetries [64, 65, 120], and this structure can be extended to QCA.
Continuing to assume that a QCA on $M$ can be identified by a map from $M$ to $\mathrm{QCA}_{d}$
$V\in\mathrm{QCA}_{M}\longleftrightarrow v:M\to\mathrm{QCA}_{d},$ (222)
lattice symmetries can be treated as actions of the symmetry group $G$ ($x\mapsto g\cdot x$) on the base manifold $M$ [120]. Then, identifying the QCA $V$ with the map $v$, the QCA is symmetric if